OpenStudy (anonymous):

Find and simplify f(-x)-f(x) for given function of f. f(x)=9x^3+2x^2+8x-11 f(-x)-f(x)=?

OpenStudy (aum):

f(x)=9x^3+2x^2+8x-11 f(-x) = ? Replace every instance of x by -x: f(-x) = 9(-x)^3 + 2(-x)^2 + 8(-x) - 11 f(-x) = -9x^3 + 2x^2 - 8x - 11 f(x) = 9x^3 + 2x^2 + 8x - 11 subtract:

OpenStudy (anonymous):

subtract what?

OpenStudy (freckles):

f(-x)-f(x)

OpenStudy (freckles):

You are to take away the polynomial labeled f(x) from the polynomial labeled f(-x)

OpenStudy (freckles):

Aum are actually found f(-x) for you The subtraction is left to you.

OpenStudy (anonymous):

im confused

OpenStudy (freckles):

f(-x) is -9x^3+2x^2-9x+11 f(x) is 9x^3+2x^2+8x+11

OpenStudy (freckles):

You are to find f(-x)-f(x)

OpenStudy (anonymous):

so put the two together?

OpenStudy (freckles):

(-9x^3+2x^2-9x+11)-(9x^3+2x^2+8x+11)

OpenStudy (freckles):

You do the subtraction that it tells you

OpenStudy (anonymous):

so just x is the answer?

OpenStudy (freckles):

(-9x^3+2x^2-8x+11)-(9x^3+2x^2+8x+11) ... so -9-9=? 2-2=? -8-8=? 11-11=?

OpenStudy (anonymous):

is it +11 for both of them? or -11?

OpenStudy (anonymous):

+11. got the answer. thanks!